# STRAWBERRY: Finding haloes in the gravitational potential

### v1

@@ 4.2 Matter density profiles
Given the previous observations, let us now investigate in more detail how the bound population is distributed by studying the matter density profile of the resulting structures. Through simulations, it has been shown that haloes on exhibit the same density profile over at least 20 orders of magnitude in halo mass (Wang et al., 2020). In particular, it has been shown that the density profile is well described by a Navarro-Frenk-White (NFW) profile (Navarro et al., 1997), which depends solely on a single scale radius and amplitude. This profile, however, has the significant shortcoming of not having a convergent mass, meaning that integrating the profile to infinitely large radii does not converge to a finite value. Because of this ambiguity over where haloes ‘end’, it has become common practice to define mass with respect to an average enclosed overdensity $\Delta$, which has the advantage of being independent of the specific profile of the halo, thus standardising the definition of ‘halo mass’. In recent years, this standardisation has been put into question. Indeed, using dynamical information (Adhikari et al., 2014), it has been shown that haloes, contrary to the most popular profiles used to describe them, have a finite extent. With several approaches recovering this feature (e.g. Diemer, 2017, 2022; Garc´ıa et al., 2023; Salazar et al., 2024).
@@
finding a good agreement between the parametric form and the density profile of the bound populations. Moreover, we find that the resulting best fit parameters are consistent with the ranges quoted in Diemer (2025), indicating that it is possible to distinguish between the orbiting and infalling populations, as defined in these works, using purely instantaneous information, meaning without the need to reconstruct particle orbits. For completeness, in App. B we fit this model to the density profiles of individual haloes and present resulting best fit parameters.
@@ 4.3 Virialisation
$2T-G=\langle{\bf v}\cdot{\bf v}\rangle-\langle{\bf r}\cdot{\bf a}\rangle=0,$ (32)
@@
In Stücker et al. (2021), the authors show that the bound population at $z=0$ is virialised with the resulting virial ratio being in agreement with the theoretical prediction. Here, we expand on these previous results and also show the redshift evolution. Indeed, the strawberry selected particles, shown in purple, yield a median in agreement with the theoretical expectation. While one may observe a slight decrease at high masses the full distribution remains consistent with the theoretical expectation. For the purpose of visual clarity, we have not shown the dispersion around the median which is roughly constant of the order of 10%. Furthermore, if we plot the median virial ratios estimated only using the particles considered unbound by strawberry (green lines in Fig. 9) we observe a very strong bias toward lower values, and also recover the redshift trend seen in the subfind selected particles. This confirms that it is indeed the inclusion of unbound particles which produces the decrease seen in classical halo finders.
@@ 4.4 Bound masses
In this final section, we investigate how our previously introduced mass definition, $M_{\rm bound}$, relates to the wider ecosystem of masses used in cosmology. Indeed, Stücker et al. (2021), observe that, at $z=0$, the bound mass appears to closely track the SO mass $M_{\rm 200b}$. Here, we wish to see if this is also the case at higher redshift. To begin, we estimate the median mass ratio $M_{\rm 200b}/M_{\rm bound}$ as a function of $M_{\rm bound}$. This ratio, shown as grey lines in Fig. 10, is computed for four redshifts, z=0, 0.5, 1.0,z=0,\ 0.5,\ 1.0, and $2.0$, corresponding with the shade of each line, respectively from brightest to darkest. We plot the resulting curves omitting the scatter around these measurements for visual clarity, the latter being roughly constant, $\sigma\simeq 0.2$, for all masses and redshift.
@@
At redshift $z=0$ we recover that the bound mass is comparable to $M_{\rm 200b}$. On the other hand, we observe a strong redshift dependence, with high redshift haloes of a given mass having larger values of $M_{\rm 200b}$ relative to $M_{\rm bound}$ than their low redshift counterparts. Simultaneously, we see a mass dependent trend emerge, indicating that, for $z>0$, lower mass haloes contain more unbound material within $r_{\rm 200b}$ than their high mass counterparts. By plotting the median mass ratio of the bound mass contained within $r_{\rm 200b}$ to the total bound mass, in pink using the same brightness convention as previously, we find that this trend is strongly reduced, indicating that for most redshifts $r_{200\rm b}$ contains the majority of bound particles.
To further understand the origin of this trend, we plot the median fraction of unbound material within $r_{200\rm b}$, in green, following the same convention for higher redshifts. In this case the redshift and mass dependence appears clearly. This indicates that while $M_{\rm 200b}$ will typically encompass the majority of the bound mass, the additional fraction of unbound material will vary with both redshift and mass. A possible explanation for the observed trend is that $r_{\rm 200b}$ is defined with respect to the total matter density and cosmological background. As such it will depend on how this matter is distributed as well as the cosmological background evolution. This creates a compounding effect where mass dependences in the density profile, for example as quantified by the concentration mass relation, and cosmological background both change the definition of $r_{200\rm b}$ (see e.g. Diemer & Kravtsov, 2014).
@@ 5 Discussion
Similarly to previous studies (e.g. Wang et al., 2011), we observe that newly accreted particles are deposited on large orbits in the outskirts of the halo. In addition, we find that the bound population is virialised and exhibits a well defined edge. On the other hand, the unbound population, which inhabits the outskirts, is not virialised and does not exhibit such an edge and even obscures the edge of the bound population when looking at the density profile of all particles surrounding the halo. In phase space and energy space, we find that this edge separates the bound population form the infalling and first splashback streams, qualitatively corresponding to a feature previously used to define the orbiting population (Garc´ıa et al., 2023; Salazar et al., 2024).
@@
This work paves the way for deeper studies of dark matter haloes. In particular, in cases where a clear definition in terms of energies is advantageous. For instance, when creating idealised initial conditions for haloes, where the relation between the density and gravitational potential is used to perform the Eddington inversion and obtain the particle distribution function (e.g. Binney & Tremaine, 2008). Alternatively, the definition of haloes outlined here can be of interest when a distinction between a virialised and non-virialised state is needed. This can be the case when studying how gas settles in the potential well, or how tidal stripping affects haloes. Finally, Ondaro-Mallea et al. (2022) associate deviations from universality in the halo mass function defined at $M_{\rm 200b}$ to the mass accretion history of haloes, given the stability of the bound profile seen in Sect. 4.2, it would be of interest to investigate further whether $M_{\rm bound}$ provides a more universal alternative to standard spherical overdensity masses.
@@ Acknowledgements.
The authors would like to thank Eduardo Rozo, Edgar Salazar, and Lurdes Ondaro-Mallea for their insightful comments. This project has received funding from the Spanish Government’s grant program “Proyectos de Generación de Conocimiento” under grant number PID2021-128338NB-I00. JS acknowledges funding by the Austrian Science Fund (FWF) [10.55776/ESP705]. Many of the numerical computations carried out, and plots presented in this work were made possible by the SciPy (Virtanen et al., 2020), NumPy (Harris et al., 2020), Matplotlib (Hunter, 2007), and Bacco (Angulo et al., 2021) packages.
@@ Appendix B Profile fitting parameters
As stated in the main text we find that, on average, the Diemer (2023) model B profile provides a good fit to the density profile of the bound population. In this appendix we provide an expanded view where we fit this model to the individual log profiles of haloes using a standard $\chi^{2}$ minimisation procedure.
@@
$r_{\rm s}$ $=(0.05\pm 0.02)\left[M_{\rm bound}/10^{13}h^{-1}\mathrm{M}_{\odot}\right]^{0.42\pm 0.01}\,h^{-1}{\rm Mpc},$ (33)
$r_{\rm t}$ $=(0.46\pm 0.01)\left[M_{\rm bound}/10^{13}h^{-1}\mathrm{M}_{\odot}\right]^{0.369\pm 0.006}\,h^{-1}{\rm Mpc},$ (34)
@@
$\alpha$ $=(0.196\pm 0.008)\left[M_{\rm bound}/10^{13}h^{-1}\mathrm{M}_{\odot}\right]^{-0.026\pm 0.004}\;\text{ and}$ (35)
$\beta$ $=(4.42\pm 0.04)\left[M_{\rm bound}/10^{13}h^{-1}\mathrm{M}_{\odot}\right]^{-0.04\pm 0.02},$ (36)

### v2

@@ 4.2 Matter density profiles
Given the previous observations, let us now investigate in more detail how the bound population is distributed by studying the matter density profile of the resulting structures. Through simulations, it has been shown that haloes on exhibit the same density profile over at least 20 orders of magnitude in halo mass (Wang et al., 2020). In particular, it has been shown that the density profile is well described by a Navarro-Frenk-White (NFW) profile (Navarro et al., 1997), which depends solely on a single scale radius and amplitude. This profile, however, has the significant shortcoming of not having a convergent mass, meaning that integrating the profile to infinitely large radii does not converge to a finite value. Because of this ambiguity over where haloes ‘end’, it has become common practice to define mass with respect to an average enclosed overdensity $\Delta$, which has the advantage of being independent of the specific profile of the halo, thus standardising the definition of ‘halo mass’. In recent years, this standardisation has been put into question. Indeed, using dynamical information (Adhikari et al., 2014), it has been shown that haloes, contrary to the most popular profiles used to describe them, have a finite extent, as shown by e.g. Diemer (2017, 2022), García et al. (2023), and Salazar et al. (2025).
@@
finding a good agreement between the parametric form and the density profile of the bound populations. Moreover, we find that the resulting best fit parameters are consistent with the ranges quoted in Diemer (2025), indicating that it is possible to distinguish between the orbiting and infalling populations, as defined in these works, using purely instantaneous information, meaning without the need to reconstruct particle orbits. For completeness, in App. B we fit this model to the density profiles of individual haloes and present the resulting best fit parameters.
@@ 4.3 Virialisation
$2T+G=\langle{\bf v}\cdot{\bf v}\rangle+\langle{\bf r}\cdot{\bf a}\rangle=0,$ (32)
@@
In Stücker et al. (2021), the authors show that the bound population at $z=0$ is virialised with the resulting virial ratio being in agreement with the theoretical prediction. Here, we expand on these previous results and also show the redshift evolution. Indeed, the strawberry selected particles, shown in purple, yield a median in agreement with the theoretical expectation. We note a residual shift towards lower values at high masses and high redshifts which we attribute to recently formed, accretion dominated haloes which have not had time to fully relax. For the purpose of visual clarity, we have not shown the dispersion around the median, which is roughly constant of the order of 10%. Furthermore, if we plot the median virial ratios estimated only using the particles considered unbound by strawberry (green lines in Fig. 9) we observe a very strong bias toward lower values, and also recover the redshift trend seen in the subfind selected particles. This confirms that it is indeed the inclusion of unbound particles which produces the decrease seen in classical halo finders.
@@ 4.4 Bound masses
In this final section, we investigate how our previously introduced mass definition, $M_{\rm bound}$, relates to the wider ecosystem of masses used in cosmology. Indeed, Stücker et al. (2021), observe that, at $z=0$, the bound mass appears to closely track the SO mass $M_{\rm 200b}$. Here, we wish to see if this is also the case at higher redshift. To begin, we estimate the median mass ratio $M_{\rm 200b}/M_{\rm bound}$ as a function of $M_{\rm 200b}$. This ratio, shown as grey lines in Fig. 10, is computed for four redshifts, z=0, 0.5, 1.0,z=0,\ 0.5,\ 1.0, and $2.0$, corresponding with the shade of each line, respectively from brightest to darkest. We plot the resulting curves omitting the scatter around these measurements for visual clarity, the latter being roughly constant, $\sigma\simeq 0.2$, for all masses and redshift.
@@
At redshift $z=0$, we recover that the bound mass is comparable to $M_{\rm 200b}$. At higher redshifts, we see that while both masses remain qualitatively similar, the ratio does shift to a lower value. This trend is driven by the bound region being more extended at higher redshifts, which can also be seen in Fig. 8. We can further observe this trend by plotting the ratio of the amount of bound mass within $r_{\rm 200b}$ with respect to the total bound mass, in pink, using the same brightness convention as previously. In doing so, we can clearly observe that this ratio increases with redshift. This indicates that at higher redshifts $r_{\rm 200b}$ contains less bound material than at lower redshift, reinforcing the observation that the bound region is more extended at higher redshifts.
To further understand the origin of this trend, we plot the median fraction of unbound material within $r_{200\rm b}$, in green, following the same convention for higher redshifts. In this case, the redshift dependence appears to be less visible. This indicates that while $M_{\rm 200b}$ will primarily contain bound material, the physical region it will be probing will change as a function of redshift. A possible explanation for the observed trend is that $r_{\rm 200b}$ is defined with respect to the total matter density and cosmological background. As such, it will not only depend on how matter is distributed inside the halo but also on the cosmological background evolution. This creates a compounding effect where changes in the density profile, as seen in Fig. 8, and cosmological background both change the definition of $r_{200\rm b}$ (see e.g. Diemer and Kravtsov, 2014).
@@ 5 Discussion
Similarly to previous studies (e.g. Wang et al., 2011), we observe that newly accreted particles are deposited on large orbits in the outskirts of the halo. In addition, we find that the bound population is virialised and exhibits a well defined edge. On the other hand, the unbound population, which inhabits the outskirts, is not virialised and does not exhibit such an edge and even obscures the edge of the bound population when looking at the density profile of all particles surrounding the halo. In phase space and energy space, we find that this edge separates the bound population form the infalling and first splashback streams, qualitatively corresponding to a feature previously used to define the orbiting population (García et al., 2023; Salazar et al., 2025).
@@
This work paves the way for deeper studies of dark matter haloes. In particular, in cases where a clear definition in terms of energies is advantageous. For instance, when creating idealised initial conditions for haloes, where the relation between the density and gravitational potential is used to perform the Eddington inversion and obtain the particle distribution function (e.g. Binney and Tremaine, 2008). Alternatively, the definition of haloes outlined here can be of interest when a distinction between a virialised and non-virialised state is needed. This can be the case when studying how gas settles in the potential well, or how tidal stripping affects haloes. Finally, Ondaro-Mallea et al. (2022) associate deviations from universality in the halo mass function defined at $M_{\rm 200b}$ to the mass accretion history of haloes, given the stability of the bound profile seen in Sect. 4.2, it would be of interest to investigate further whether $M_{\rm bound}$ provides a more universal alternative to standard spherical overdensity masses.
@@ Acknowledgements.
The authors would like to thank Eduardo Rozo, Edgar Salazar, Lurdes Ondaro-Mallea, and the anonymous referee for their insightful comments. This project has received funding from the Spanish Government’s grant program “Proyectos de Generación de Conocimiento” under grant number PID2021-128338NB-I00. JS acknowledges funding by the Austrian Science Fund (FWF) [10.55776/ESP705]. Many of the numerical computations carried out, and plots presented in this work were made possible by the SciPy (Virtanen et al., 2020), NumPy (Harris et al., 2020), Matplotlib (Hunter, 2007), and Bacco (Angulo et al., 2021) packages.
@@ Appendix B Profile fitting parameters
As stated in the main text we find that, on average, the Diemer (2023) model B profile provides a good fit to the density profile of the bound population. In this appendix, we provide an expanded view where we fit this model to the individual log profiles of haloes using a standard $\chi^{2}$ minimisation procedure.
@@
$r_{\rm s}$ $=(0.05\pm 0.01)\left[M_{\rm bound}/10^{13}h^{-1}\mathrm{M}_{\odot}\right]^{0.403\pm 0.007}\,h^{-1}{\rm Mpc},$ (33)
$r_{\rm t}$ $=(0.46\pm 0.01)\left[M_{\rm bound}/10^{13}h^{-1}\mathrm{M}_{\odot}\right]^{0.352\pm 0.007}\,h^{-1}{\rm Mpc},$ (34)
@@
$\alpha$ $=(0.20\pm 0.01)\left[M_{\rm bound}/10^{13}h^{-1}\mathrm{M}_{\odot}\right]^{-0.041\pm 0.008}\;\text{ and}$ (35)
$\beta$ $=(4.15\pm 0.03)\left[M_{\rm bound}/10^{13}h^{-1}\mathrm{M}_{\odot}\right]^{-0.05\pm 0.01},$ (36)